# Addressing Model and Data Heterogeneity in Multimodal Large Language Model Training

### v1

@@ 5 Addressing Data Heterogeneity
The data heterogeneity causes intra-microbatch and inter-microbatch stragglers. Addressing both straggler simultaneously renders the problem intractable due to its high complexity (e.g., the order of the training samples results in factorial-level complexity) and the necessity of runtime execution. To tackle this challenge, we decompose the overarching problem into two tractable subproblems. Initially, we present intra-microbatch reordering to eliminate stragglers across DP groups. We then introduce inter-microbatch reordering to minimize pipeline bubbles caused by the straggler microbatches. The combination of the two reordering, i.e., data-aware reordering, effectively addresses the data heterogeneity.
@@ 5.1 Intra-microbatch Reordering
Insight. To address the first subproblem of intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples. As illustrated in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples. To neutralize this imbalance, we propose reordering the training samples within the global batch by size. Specifically, as depicted in Figure 11, we reorder the training samples into the sequence . This strategy effectively distributes the computational load more evenly and improves the scalability. It eliminates the straggler without breaking the synchronous training semantics [31], as the reordering occurs within an individual global batch.
Intra-microbatch Reordering. Leveraging this insight, we propose intra-microbatch reordering to balance computational load and improve the overall scalability. Formally, the challenge involves minimizing the maximum computation time among DP groups. This problem corresponds to the NP-hard problem, multiway number partitioning [36], which aims to minimize the largest sum. There is no known polynomial-time algorithm capable of determining the optimal solution. Given the variability of runtime training samples, any algorithm employed must run before each iteration with negligible runtime overhead. Consequently, we adopt the greedy number partitioning algorithm, which guarantees an approximation ratio of [37]. The detailed algorithm is summarized in Algorithm 2. The function receives the original training samples and DP size . This algorithm first sorts the training samples in ascending order by the sample size (line 3). Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8). It then returns the reordered training samples (line 9-11). The algorithm is lightweight and maintains a time complexity of .
@@ 5.2 Inter-microbatch Reordering
The second subproblem is the inter-microbatch straggler. As we discussed in §2.3, data heterogeneity leads to varied computation times across microbatches within the modality encoder and generator. The straggler microbatch prolongs the training by creating large pipeline bubbles. In the context of 1F1B pipeline scheme, the overall iteration time is primarily governed by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12. Let the PP size be and the number of microbatches be ( and in Figure 12). Typically, is larger than to reduce the proportion of time spent in the warm-up and cool-down phases. We abstract an innovative concept, pipeline intervals, at the first PP stage. As shown in Figure 12, these intervals are typically filled with forward pass, except for the last intervals (i.e., , , and ). Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble). Designing an efficient runtime algorithm to minimize pipeline bubbles presents a significant challenge.
Insights. We leverage two insights to solve this problem. The first insight involves minimizing the volume of intervals that cannot be filled. As shown in Figure 12, the last intervals (i.e., to ) remain unfilled. These intervals become the pipeline bubbles and prolong the training iteration. We observe a positive correlation between the volume of and the size of the microbatch. For instance, is significantly larger than and . By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals (i.e., pipeline bubbles).
The second insight involves minimizing the unfilled area of front intervals. As shown in Figure 12, the first interval is filled with by the to microbatches’ forward passes. The volume of the first interval (i.e., ) is determined by the aggregate forward and backward times of the first microbatch in LLM backbone and modality generator. Such characteristic allows the accommodation of large microbatches early in the sequence to fill the first interval. For subsequent intervals, is filled by the microbatch’s forward pass. By evaluating the volume of , we place the microbatch, whose forward time most closely matches this volume, at the position, thereby reducing the unfilled area (i.e., pipeline bubbles).
@@
The functions and operate with a time complexity of . The function calculates interval volumes using the current order . This calculation is facilitated by a dynamic programming algorithm that utilizes a recursive formula derived from pipeline dependencies. Specifically, each microbatch’s start time depends on two factors: the completion of the preceding microbatch on the same device and the availability of input data from the upstream microbatch. Consequently, the end time of each microbatch is determined by the maximum of these two dependencies plus its own computation time. This dynamic programming algorithm exhibits a complexity of per function invocation. The algorithm is lightweight and efficient with a time complexity of .
@@ 6 System Implementation and Optimization
We implement MMScale with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models. MMScale is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §7). MMScale leverages a distributed file system to store the training data and model checkpoints. It handles failures by automatically recovering the training process from the latest checkpoint.
MMScale orchestrator. MMScale’s training orchestrator, implemented as a Python script on a dedicated CPU node, formulates the adaptive resource allocation problem using Disciplined Convex Programming [38]. It employs the CVX solver [39] to efficiently solve this problem within milliseconds. The orchestrator then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
MMScale initializer. MMScale incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs. Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the default communication backend, except for the TP communication where we use StepCCL instead. MMScale shards the model parameters in accordance with the established parallelism strategy.
Parallelism unit. As we discussed in §4.1, the disaggregated training is implemented through a specific module: parallelism unit. When initializing the distributed training, MMScale first establishes the communication groups within one parallelism unit. Each GPU process has a global rank and a local rank within the unit, which facilitates the distributed initialization. Then, it initializes the communication broker to establish the PP communication between adjacent parallelism units. All communication traffic between parallelism units is routed via the communication broker.
@@
Asynchronous data loader. When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data. The latter significantly increases data loading time. For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images. The text is just kilobytes, whereas the images are total of 120 megabytes. Loading multiple such samples can take several seconds. MMScale’s data loader employs a producer-consumer model to enhance efficiency. The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and pre-processes training data (e.g., image cropping) asynchronously with the GPU training process. The consumer, integrated within the main training process, receives this pre-processed data and reorders it to eliminate the data heterogeneity. The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available.
@@ 7 Evaluation
In this section, we first use large-scale experiments to demonstrate the overall performance improvements of MMScale over existing multimodal LLM training solution. Next, we use microbenchmarks to deep dive into MMScale and show the effectiveness of each component of MMScale. Finally, we conduct a case study to evaluate MMScale under different frozen settings and evaluate the system optimizations.
@@
Models. For LLM backbone, we choose the representative LLM architecture, Llama3 [22], which is widely used in both academia and industry. Table 1 lists the detailed model configurations. As for the modality, we focus on images and texts. MMScale is also compatible with other modalities. For modality encoder and modality generator, we use ViT-Huge [45] (0.63B) and Stable-Diffusion 2.1 [46] (1B) respectively. These two models are widely used for image understanding and generation. The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and Stable-Diffusion to form multimodal LLMs designated as MLLM-9B, MLLM-15B, and MLLM-72B. For large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information. For small models, we use low image resolution (i.e., 512512).
@@
Metrics. We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate MMScale. MFU measures the percentage of GPU FLOPs that are effectively utilized during model training. We also use the training throughput (TPT.) to evaluate the training speed of MMScale. Since MMScale and Megatron-LM may utilize different numbers of GPUs due to varying resource allocation strategies, we also indicate the number of GPUs used in the throughput charts.
@@ 7.1 Overall Performance
Setup. We first compare the overall performance of MMScale against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs). We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline. Megatron-LM employs rigid resource allocation as described in §2.1. In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B, respectively. TP size is set to 8. As for MMScale, the parallelism strategy is determined by adaptive resource allocation. In our experiments, one GPU is able to facilitate training ViT and Stable-Diffusion. For both Megatron-LM and MMScale, we replicate the encoder and generator across the GPUs within the TP group to process different image subsequences, whereas TP itself is not used. We set global batch size to 1920.
@@
The experimental results are shown in Figure 15 and Figure 16. Figure 15 shows the MFU of MMScale and Megatron-LM under different multimodal LLMs. Figure 16 shows the training throughput and marks the number of GPUs used in each experiment. We summarize the experiment as follows.
@@
As shown in Figure 15, MMScale achieves 51.8%-54.7% MFU under large-scale multimodal LLM training scenario. This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of MMScale in eliminating the model and data heterogeneity in multimodal LLM training.
@@
MMScale significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs. These performance gains largely stem from MMScale’s adaptive resource allocation strategy. Megatron-LM’s rigid strategy often leads to GPU underutilization in multimodal modules, since it assigns too many GPUs to the modality encoder and generator. In contrast, MMScale adaptively adjusts GPU allocations based on specific model and data demands. Additionally, MMScale’s data-aware reordering technique further boosts efficiency.
@@
In MLLM-72B training scenario, MMScale also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs. The high image resolution prolongs the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone. MMScale addresses this by allocating additional GPUs to these modules to balance the pipeline. MMScale’s data-aware reordering strategy continues to diminish data heterogeneity, thereby increasing overall training efficiency.
7.2 Deep Dive into MMScale
@@
In this subsection, we perform microbenchmarks to evaluate the effectiveness of each MMScale’s component. Our experiments utilize up to 98 NVIDIA A800 GPUs. We set the global batch size to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
7.2.1 Adaptive Resource Allocation
@@
We measure the MFU and training throughput of MMScale and other resource allocation strategies, including a comparison with Megatron-LM and a ratio-based approach. The ratio-based strategy allocates GPUs according to the computational demands (flops) of each module. We also annotate the total number of GPUs utilized in the throughput chart. The experimental results are shown in Figure 17. MMScale consistently outperforms the baseline strategies, achieving 1.3%-2.7% higher MFU and 1.4%-2.7% higher training throughput. Although the ratio-based strategy outperforms Megatron-LM’s rigid allocation, it still lags behind MMScale since it neglects the intricate performance model (§4.2) of multimodal LLM training. MMScale employs an adaptive resource allocation strategy that optimally balances computational loads across the three modules and achieves high resource utilization.
We also evaluate the running time of MMScale’s resource allocation algorithm under different training settings, as detailed in Table 2. The algorithm completes the allocation process in under one second. The overhead is negligible compared to the days or even weeks required for overall training.
7.2.2 Data-aware Reordering
We evaluate the effectiveness of MMScale’s data-aware reordering by comparing it against the random order set by PyTorch’s DataLoader, while keeping other components the same. The effectiveness is gauged through metrics such as MFU and training throughput (TPT.). We use the optimal resource allocation and parallelism strategy decided by MMScale’s adaptive resource allocation algorithm. Given that the GPU allocation policy remains unchanged, the number of GPUs is not specified in this experiment. The experimental settings are the same as those in §7.2.1. The results are shown in Figure 18. MMScale consistently outperforms the baseline, achieving 1.03%-1.11% higher MFU and training throughput. The performance gap becomes more pronounced as the model size decreases. This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more inter-microbatch heterogeneity. In essence, MMScale’s data-aware reordering effectively mitigates data heterogeneity and enhances the training efficiency.

### v2

@@ 6 Addressing Data Heterogeneity
Disaggregated training enables preprocessing the training data with negligible runtime overhead. Based on this, we introduce disaggregated data reordering, seamlessly integrated into data preprocessing, to address data heterogeneity without additional overhead. Initially, we present intra-microbatch reordering to eliminate stragglers across DP groups. We then introduce inter-microbatch reordering to minimize pipeline bubbles caused by the straggler microbatches. The combination of the two reordering, i.e., disaggregated data reordering, effectively addresses the data heterogeneity. We emphasize that these two reordering algorithms are integrated into the disaggregated data preprocessing. This ensures that the complex reordering does not interfere with the GPU training process.
@@ 6.1 Intra-microbatch Reordering
Insight. To address the first subproblem, intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples. As illustrated in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples. To neutralize this imbalance, we propose reordering the training samples within the global batch by size. Specifically, as depicted in Figure 11, we reorder the training samples into the sequence . This strategy effectively distributes the computational load more evenly and improves the scalability. It eliminates the straggler without breaking the synchronous training semantics [31], as the reordering occurs within an individual global batch.
Intra-microbatch Reordering. Leveraging this insight, we propose intra-microbatch reordering to balance computational load and improve the overall scalability. Formally, the challenge involves minimizing the maximum computation time among DP groups. This problem corresponds to the NP-hard problem, multiway number partitioning [36], which aims to minimize the largest sum. There is no known polynomial-time algorithm capable of determining the optimal solution. Given the substantial batch size in production training, the algorithm employed must be lightweight and efficient. Consequently, we adopt the greedy number partitioning algorithm, which guarantees an approximation ratio of [37]. The detailed algorithm is summarized in Algorithm 2. The function receives the original training samples and DP size . This algorithm first sorts the training samples in ascending order by the sample size (line 3). Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8). It then returns the reordered samples (line 9-11). The algorithm has a time complexity of .
@@ 6.2 Inter-microbatch Reordering
The second subproblem is the inter-microbatch straggler. As we discussed in §2.3, data heterogeneity leads to varied computation times across microbatches within the modality encoder and generator. The straggler microbatch prolongs the training by creating large pipeline bubbles. In the context of 1F1B pipeline scheme, the overall iteration time is primarily governed by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12. Let the PP size be and the number of microbatches be ( and in Figure 12). Typically, is larger than to reduce the proportion of time spent in the warm-up and cool-down phases. We abstract an innovative concept, pipeline intervals, at the first PP stage. As shown in Figure 12, these intervals are typically filled with forward pass, except for the last intervals (i.e., , , and ). Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble).
Insights. We leverage two insights to solve this problem. The first insight involves minimizing the volume of intervals that are not filled. As shown in Figure 12, the last intervals (i.e., to ) remain unfilled. These intervals become the pipeline bubbles and prolong the training iteration. We observe a positive correlation between the volume of and the size of the microbatch. The size refers to the computation time of the microbatch in modality encoder and generator. For instance, is significantly larger than and since the microbatch is the largest. By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals (i.e., pipeline bubbles).
The second insight involves minimizing the unfilled area of left intervals. The left intervals (i.e., to ) are filled with the forward pass. As shown in Figure 12, the first interval is filled with by the to forward passes. For subsequent intervals, is filled by the forward pass. By evaluating the volume of , we place the microbatches, whose forward time most closely matches this volume, at the corresponding position, to minimize the unfilled area (i.e., pipeline bubbles).
@@
The functions and operate with a time complexity of . The function calculates interval volumes using the current order . This calculation is facilitated by a dynamic programming algorithm that utilizes a recursive formula derived from pipeline dependencies. Specifically, each microbatch’s start time depends on two factors: the completion of the preceding microbatch on the same device and the availability of input data from the upstream microbatch. Consequently, the end time of each microbatch is determined by the maximum of these two dependencies plus its own computation time. This dynamic programming algorithm exhibits a complexity of per function invocation. The algorithm has a time complexity of .
@@ 7 System Implementation and Optimization
We implement DistTrain with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models. DistTrain is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §8). DistTrain leverages a distributed file system to store the training data and model checkpoints. It handles failures by automatically recovering the training process from the latest checkpoint.
DistTrain manager. DistTrain’s training manager, implemented as a Python script on a dedicated CPU node, formulates the disaggregated model orchestration problem using Disciplined Convex Programming [38]. It employs the CVX solver [39] to efficiently solve this problem within milliseconds. The manager then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
DistTrain initializer. DistTrain incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs. Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the communication backend, except for the TP communication where we use StepCCL instead. DistTrain shards the model parameters in accordance with the established parallelism strategy.
Parallelism unit. As we discussed in §4.1, the GPU training disaggregation is implemented through a specific module: parallelism unit. When initializing the distributed training, DistTrain first establishes the communication groups within one parallelism unit. Each GPU process has a global rank and a local rank within the unit, which facilitates the distributed initialization. Then, it initializes the communication broker to establish the PP communication between adjacent parallelism units. All communication traffic between parallelism units is routed via the communication broker.
@@ 8 Evaluation
In this section, we first use large-scale experiments to demonstrate the overall performance improvements of DistTrain over Megatron-LM. Next, we use microbenchmarks to deep dive into DistTrain and show the effectiveness of each component in DistTrain. Finally, we provide a case study to further elucidate DistTrain’s capabilities.
@@
Models. For LLM backbone, we choose the representative LLM architecture, Llama3 [22], which is widely used in both academia and industry. Table 1 lists the detailed model configurations. As for the modality, we focus on images and texts. DistTrain is also compatible with other modalities. For modality encoder and modality generator, we use ViT-Huge [45] (0.63B) and Stable-Diffusion 2.1 [46] (1B) respectively. These two models are widely used for image understanding and generation. The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and Stable-Diffusion to form multimodal LLMs designated as MLLM-9B, MLLM-15B, and MLLM-72B. For large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information. For small models, we use low image resolution (i.e., 512512).
@@
Metrics. We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate DistTrain. MFU measures the percentage of GPU FLOPs that are effectively utilized during model training. We also use the training throughput (TPT.) to evaluate the training speed of DistTrain. Since DistTrain and Megatron-LM may utilize different numbers of GPUs due to varying model orchestration strategies, we also indicate the number of GPUs used in the throughput charts.
@@ 8.1 Overall Performance
Setup. We first compare the overall performance of DistTrain against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs). We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline. Megatron-LM employs rigid model orchestration as described in §2.1. In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B. PP size is set to 1 for modality encoder and generator. TP size is set to 8. As for DistTrain, the parallelism strategy is determined by disaggregated model orchestration. In our experiments, one GPU is able to facilitate training ViT and Stable-Diffusion. We replicate the encoder and generator across the GPUs within the TP group to process different images, whereas TP itself is not used. We set global batch size to 1920.
@@
The experimental results are shown in Figure 15 and Figure 16. Figure 15 shows the MFU. Figure 16 shows the training throughput and marks the number of GPUs used in each experiment. The different GPU numbers are due to the varying model orchestration strategies of DistTrain and Megatron-LM. We summarize the experiment as follows.
@@
As shown in Figure 15, DistTrain achieves 51.8%-54.7% MFU in large-scale multimodal LLM training. This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of DistTrain in addressing the model and data heterogeneity in multimodal LLM training.
@@
DistTrain significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs. These performance gains largely stem from DistTrain’s disaggregated model orchestration. Megatron-LM’s rigid strategy often leads to GPU underutilization, since it assigns too many GPUs to the modality encoder and generator. In contrast, DistTrain adaptively adjusts model orchestration based on specific model and data demands. Additionally, DistTrain’s disaggregated data reordering technique further boosts efficiency.
@@
In MLLM-72B training scenario, DistTrain also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs. The high image resolution prolongs the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone. DistTrain addresses this by allocating additional GPUs to these modules to balance the pipeline. The disaggregated data reordering strategy continues to diminish data heterogeneity, thereby increasing training efficiency.
8.2 Deep Dive into DistTrain
@@
In this subsection, we perform microbenchmarks to evaluate the effectiveness of each DistTrain’s component. Our experiments utilize up to 98 NVIDIA A800 GPUs. We set the global batch size to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
8.2.1 Disaggregated Model Orchestration
@@
We measure the MFU and training throughput of DistTrain and other model orchestration strategies, including a comparison with Megatron-LM and a ratio-based approach. The ratio-based strategy allocates GPUs according to the computational demands (flops) of each module. We also annotate the total number of GPUs utilized in the throughput chart. The experimental results are shown in Figure 17. DistTrain consistently outperforms the baseline strategies, achieving 1.3-2.7 higher MFU and 1.4-2.7 higher training throughput. Although the ratio-based strategy outperforms Megatron-LM’s rigid strategy, it still lags behind DistTrain since it neglects the intricate performance model (§5.1) of multimodal LLM training. DistTrain’s disaggregated model orchestration optimally balances computational loads across the three modules and achieves high resource utilization.
@@
We also evaluate the running time of DistTrain’s disaggregated model orchestration under different training settings, as detailed in Table 2. The algorithm completes in under one second. The overhead is negligible compared to the days or even weeks required for overall training.
8.2.2 Disaggregated Data Reordering
We evaluate the effectiveness of DistTrain’s disaggregated data reordering by comparing it against the random order, while keeping other components the same. The effectiveness is gauged through metrics such as MFU and training throughput (TPT.). We use the optimal resource allocation and parallelism strategy decided by DistTrain’s disaggregated model orchestration. Given that the model orchestration strategy remains unchanged, the number of GPUs is not shown. The experimental settings are the same as those in §8.2.1. The results are shown in Figure 18. DistTrain consistently outperforms the baseline, achieving 1.03-1.11 higher MFU and training throughput. The performance gap becomes more pronounced as the model size decreases. This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more inter-microbatch heterogeneity. In essence, DistTrain’s disaggregated data reordering effectively mitigates data heterogeneity and enhances the training efficiency. We do not measure the running time of the reordering algorithm as it operates on dedicated CPU nodes asynchronously. It does not interfere with the GPU training process. The only overhead is the network delay which is evaluated in §8.3.